OpenStudy (anonymous):

3(4x -2) - 14 = 2(6 - 2x)

OpenStudy (kittiwitti1):

Do the Distributive Property and then isolate x... :-) 3(4x-2) = ? 2(6-2x) = ?

OpenStudy (anonymous):

12x-6-14=12-4x

OpenStudy (kittiwitti1):

O_O that was fast... Alright yes, can you do it from here? :-D Just isolate the x ^-^

OpenStudy (anonymous):

-8=8x?

OpenStudy (kittiwitti1):

Well, let's see.\[12x-6-14=12-4x\]\[12x-20=12-4x\]\[12x+4x=22\]\[16x=22\]...? *checks on paper* lol

OpenStudy (kittiwitti1):

Yep, it's right o_e Does the question have a fraction/decimal answer?

OpenStudy (anonymous):

x = -4 x = -2 x = -1 x = 2 x = 3 no the answers are

OpenStudy (kittiwitti1):

O_O ... ?!

OpenStudy (anonymous):

and i had got -1

OpenStudy (kittiwitti1):

oh? how did you do the problem :-O

OpenStudy (anonymous):

subtracted 12-6-14 i had got -8 then i subtracted 12-4 and got 8 so then i divided those and got -1

OpenStudy (kittiwitti1):

Oh... idk how I did it wrong at first but I believe the answer is 16x=32. So starting from here; 12x-6-14=12-4x...\[12x-6=26-4x\]\[16x=32...\]

OpenStudy (kittiwitti1):

It's not 12-4, it's 12-(-4), you have a sign error

OpenStudy (anonymous):

so doing it your way x=2

OpenStudy (kittiwitti1):

Yep https://www.wolframalpha.com/input/?i=12x-6-14%3D12-4x

OpenStudy (anonymous):

12x - 6 -14 = 12 - 4x 12x -20 = 12 - 4x 12x + 4x =12 + 20 16 x = 32 x =32/16 x=2

OpenStudy (kittiwitti1):

Yeah, I did that. I don't know how I got it wrong at first, but it's that lol

OpenStudy (kittiwitti1):

Proof: https://www.wolframalpha.com/input/?i=3%284x-2%29-14%3D2%286-2x%29 It's x=2 lol

OpenStudy (anonymous):

lol thank you again

OpenStudy (kittiwitti1):

OH I know what I did wrong lol I put 12+20 as 22... wow what a dumb mistake xD

OpenStudy (kittiwitti1):

No problem :-D

OpenStudy (anonymous):

do you know this one The sum of two numbers is 50. If one number is 21, the other number is 18 21 26 29 30

OpenStudy (kittiwitti1):

Well.........

OpenStudy (kittiwitti1):

What's a sum? :-P

OpenStudy (anonymous):

the total amount?

OpenStudy (kittiwitti1):

I'll give you a tip though... one number is 21 and the other number is unknown.... let's make that other number "x" If you know what a sum is... which sign you use (add, divide, multiply, subtract), then you can solve for it :-)

OpenStudy (anonymous):

add

OpenStudy (kittiwitti1):

Yes it is the total amount... so which sign do you use. Yep! :-D

OpenStudy (kittiwitti1):

So what is the equation you make from these given values? x, 21 sum is 50...

OpenStudy (kittiwitti1):

am I confusing you lol

OpenStudy (anonymous):

kind of lol

OpenStudy (kittiwitti1):

okay so sum is add two numbers... and you have the two numbers: 21 and x the sum (total) is 50... you got it now? :-P

OpenStudy (anonymous):

so it would be 29

OpenStudy (kittiwitti1):

Yes lol

OpenStudy (anonymous):

thanks